OpenStudy (anonymous):

How would I simplify √75*2√48

OpenStudy (anonymous):

√75*2√48=√3*25 * 2 * √3*16= 5*√3 * 2* 4* √3= 40 *3= 120

OpenStudy (anonymous):

so the answer would 40*3=120

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you for the break down :)